(D^2-3D+2)y=e^x/1+e^x

1
Expert's answer
2022-01-17T16:11:47-0500

Find solution of the associated homogeneous equation:


"(D^2-3D+2)y=0"

Characteristic (auxiliary) equation is


"r^2-3r+2=0"

"(r-1)(r-2)=0"

"r_1=1, r_2=2"

The general solution of the associated homogeneous equation is


"y_h=c_1e^x+c_2e^{2x}"

Variation of parameters


"y'=c_1'e^x+c_1e^x+c_2'e^{2x}+2c_2e^{2x}"

"c_1'e^x+c_2'e^{2x}=0"

"y'=c_1e^x+2c_2e^{2x}"

"y''=c_1'e^x+c_1e^x+2c_2'e^{2x}+4c_2e^{2x}"

Substitute


"c_1'e^x+c_1e^x+2c_2'e^{2x}+4c_2e^{2x}"

"-3c_1e^x-6c_2e^{2x}+2c_1e^x+2c_2e^{2x}=\dfrac{e^x}{1+e^x}"

We have


"c_1'e^x+c_2'e^{2x}=0"

"c_1'e^x+2c_2'e^{2x}=\dfrac{e^x}{1+e^x}"

Then


"c_1'=-c_2'e^x"

"c_2'e^{x}=\dfrac{1}{1+e^x}"

Or


"c_1'=-\dfrac{1}{1+e^x}"

"c_2'=\dfrac{e^{-x}}{1+e^x}"

Integrate


"c_1=-\int \dfrac{1}{1+e^x}dx=-\int \dfrac{1+e^x-e^x}{1+e^x}dx"

"=-\int dx+ \int \dfrac{e^x}{1+e^x}dx=-x+\ln(e^x+1)"

"c_2=\int \dfrac{e^{-x}}{1+e^x}dx"

"e^x=u=>du=e^xdx=>dx=\dfrac{du}{u}"


"c_2=\int \dfrac{e^{-x}}{1+e^x}dx=\int \dfrac{du}{u^2(1+u)}"

"\dfrac{1}{u^2(1+u)}=\dfrac{A}{u}+\dfrac{B}{u^2}+\dfrac{C}{1+u}"

"=\dfrac{Au(1+u)+B(1+u)+Cu^2}{u^2(1+u)}"

"u=0: B=1"

"u=-1: C=1"

"u=1:2A+2B+C=1=>A=-1"

"c_2=\int \dfrac{e^{-x}}{1+e^x}dx=\int \dfrac{du}{u^2(1+u)}"

"=-\int \dfrac{du}{u}+\int \dfrac{du}{u^2}+\int \dfrac{du}{1+u}"

"=-\ln |u|-\dfrac{1}{u}+\ln(|1+u|)"

"=-x-e^{-x}+\ln(1+e^x)"

The particular solution of the non homogeneous differential equation is


"y_p=-xe^x+e^x\ln(e^x+1)"

"-xe^{2x}-e^{x}+e^{2x}\ln(1+e^x)"

The general solution of the non homogeneous differential equation is


"y=c_1e^x+c_2e^{2x}-xe^x+e^x\ln(e^x+1)"

"-xe^{2x}-e^{x}+e^{2x}\ln(1+e^x)"